OpenStudy (anonymous):

Find Three Equations that have an x-intercept at (-7,0)

OpenStudy (shamim):

x=-7

OpenStudy (anonymous):

thts not rlly an equation

OpenStudy (anonymous):

ok one is y=-3x+21

OpenStudy (anonymous):

then y=-2+14

OpenStudy (shamim):

why not equation

OpenStudy (anonymous):

i meant y=-2x+14

OpenStudy (anonymous):

then y=-4x+28

OpenStudy (anonymous):

there u go

OpenStudy (shamim):

y=-7 is an equation of line which is parallel to y axis

OpenStudy (anonymous):

wait shakur

OpenStudy (anonymous):

change my equations i give to positive at the beginning

OpenStudy (anonymous):

y=4x+28 y=3x+21 y=2x+14

OpenStudy (anonymous):

there u go

OpenStudy (anonymous):

y=4x+28 To find the x-intercept, substitute in 0 for y and solve for x. (0)=4x+28 Since x is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. 4x+28=0 Since 28 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 28 from both sides. 4x=−28 Divide each term in the equation by 4. 4x4=−284 Simplify the left-hand side of the equation by canceling the common factors. x=−284 Simplify the equation. x=−7 To find the y-intercept, substitute in 0 for x and solve for y. y=4(0)+28 Multiply 4 by each term inside the parentheses. y=0+28 Solve the equation. y=28 These are the x and y intercepts of the equation y=4x+28. x-intercept: (−7,0) y-intercept: (0,28)

OpenStudy (anonymous):

284=28/4

OpenStudy (anonymous):

4x4=4x/4

OpenStudy (anonymous):

in my explanation

OpenStudy (anonymous):

But I have a major problem in graphing

OpenStudy (anonymous):

?

OpenStudy (anonymous):

listen the x and y intercepts i showed u how to get, u just plot them on a graph, and connect the dots

OpenStudy (anonymous):

x intercepts is when the line hits the x axis that goes from left to right

OpenStudy (anonymous):

y intercept is when the line hits the y axis that goes up adn down in a grpah

OpenStudy (anonymous):

just plot -7 and 28 on a graph

OpenStudy (anonymous):

then connect

OpenStudy (anonymous):

I will really need to work on this before my final

OpenStudy (anonymous):

Here is some question: Find the intercepts and then graph the line x+y=-2

OpenStudy (anonymous):

x-intercept: (−2,0) y-intercept: (0,−2)

OpenStudy (anonymous):

x+y=−2 To find the x-intercept, substitute in 0 for y and solve for x. x+(0)=−2 Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. x=−2 To find the y-intercept, substitute in 0 for x and solve for y. (0)+y=−2 Solve the equation. y=−2 These are the x and y intercepts of the equation x+y=−2. x-intercept: (−2,0) y-intercept: (0,−2)

OpenStudy (anonymous):

here ill explain more

OpenStudy (anonymous):

x+y=−2 To find the x-intercept, substitute in 0 for y and solve for x. x+(0)=−2 Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. Remove the parentheses around the expression 0. x+0=−2 Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. x=−2 x=−2 To find the y-intercept, substitute in 0 for x and solve for y. (0)+y=−2 Solve the equation. Remove the parentheses around the expression 0. 0+y=−2 Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. y=−2 y=−2 These are the x and y intercepts of the equation x+y=−2. x-intercept: (−2,0) y-intercept: (0,−2)

OpenStudy (anonymous):

I am trying to draw the graph and see if I will get it right

OpenStudy (anonymous):

heres the graph

OpenStudy (anonymous):

wow are you a math instructor ,how can I be good in algebra

OpenStudy (anonymous):

Find the intercepts and then graph the line -4x+2y=8